OpenStudy (alyssajobug):

Solve the equation for the values of x. |4 + 5x| + 3x = 12

OpenStudy (anonymous):

X =1

OpenStudy (jdoe0001):

\(\bf |4 + 5x| + 3x = 12\implies |4 + 5x| =12-3x\implies \begin{cases} +(4 + 5x) =12-3x\\ \quad \\ \bf -(4 + 5x) =12-3x \end{cases}\)

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf |4 + 5x| + 3x = 12\\ \quad \\\implies |4 + 5x| =12-3x\implies \begin{cases} +(4 + 5x) =12-3x\\ \quad \\ \bf -(4 + 5x) =12-3x \end{cases}\)

OpenStudy (alyssajobug):

What? I don't understand any of that...

OpenStudy (jdoe0001):

an absolute value expression, gives 2 results, because it splits into two scenarios, those 2 above so you solve both and you'd get 2 values for "x"

OpenStudy (jdoe0001):

all we did firstly is, isolate the absolute value expression on the left-hand side

OpenStudy (mathstudent55):

To solve an absolute value equation of the form \( |X| = b\) where \(X\) is an expression in x, split it up into the two equations \( X = b\) or \(-X = b\) Then solve the two equations and separate the answers with the word "or."

OpenStudy (mathstudent55):

This is what @jdoe0001 explained above. You started with |4 + 5x| + 3x = 12 The first step is to change it to |4 + 5x| = 12 - 3x by subtracting 3x from both sides. Now apply the rule above: 4 + 5x = 12 - 3x or -(4 + 5x) = 12 - 3x This is the result that jdoe001 showed above. Now you solve both equations and separate the answers with the word "or."

OpenStudy (alyssajobug):

OH! Ok, Thanks for the breakdown @mathstudent55 !!! SOOOO HELPFUL!

OpenStudy (mathstudent55):

You're welcome.